# Data-Driven Characterization of Latent Dynamics on Quantum Testbeds

### v1

@@ 3.3 Training Procedure
The training procedure for learning the structure preserving model as well as the neural network-based source term amounts to solving an optimization that is constrained by the augmented dynamical equation (4). Denote by $\boldsymbol{\xi}$ a set of experiments parameterized by control functions that determine the time-evolution of the true/exact quantum state. For the numerical examples below, this set consists of various microwave control pulses, $p(t),q(t)$ in (2) as described in Section 4. Each experiment defines the (true/exact) evolution of the density matrix, $\widetilde{\rho}(t_{j};\xi_{i})$, which is estimated using quantum state tomography [43] at various time steps $t_{j}\in(0,T]$. In this work, we apply linear inversion followed by spectral filtering and renormalization, Eq. 11. Let $\boldsymbol{\theta}$ denote the set of learnable parameters in the augmented differential equation, i.e. $\boldsymbol{\theta}=\left\{\alpha_{l}\bigcup\gamma_{l}:l=1,\ldots,N^{2}-1\right\}$ for the structure preserving ansatz in Section 3.1 and $\boldsymbol{\theta}=\left\{\mathbf{W}_{l}\bigcup\mathbf{b}_{l}:l=1,\ldots,L\right\}$ for the neural network model defined in Section 3.2. The goal of the training procedure is then to find optimal parameters $\boldsymbol{\theta}$ such that the resulting augmented differential equation matches the measured data $\widetilde{\rho}$ at time-steps $t_{j}$. To this end, the following constraint optimization problem is solved:
@@
In order to evaluate the objective function, the augmented dynamical system (12b) is evolved numerically for the given experimental setup in $\boldsymbol{\xi}$ and the current parameters $\theta$. We here use an explicit 4th-order Runge-Kutta scheme for the numerical time-integration of the dynamical system, although any other numerical solver can be used. This yields the predicted evolution $\rho(t_{j})$ which is compared to the true evolution within the loss function in terms of their Frobenius distance. We utilize the Julia SciML [44] framework to solve this optimization problem with the ADAM optimizer mini-batches as an initial optimizer, followed by a full-batch L-BFGS optimization. The gradient is calculated using automatic differentiation.
After training on a defined set of experiments and time-horizon $0<t_{j}<T$, we investigate the efficacy of the trained augmented model on a set of validation experiments. In particular, we show that the trained model is able to predict the time-evolution on time domains much longer than the training domain, for experimental setups that have not been included in training. We refer to the accuracy over the training and validation sets as interpolation and extrapolation accuracy, respectively.
In the sections that follow, we will compare the solutions obtained using different formulations using the trace distance $T(\widetilde{\rho},\rho)=\frac{1}{2}\mathrm{Tr}\left[\sqrt{(\widetilde{\rho}-\rho)^{\dagger}(\widetilde{\rho}-\rho)}\right]$. In each case, we will apply the spectral filter in Eq. 11 to obtain a valid density matrix. To quantify the generalization (extrapolation) properties of the trained models, we employ the expected trace distance
$\mathbb{E}\left[T\left(\widetilde{\rho},\rho(\boldsymbol{\theta})\right)\right]=\int_{\Omega_{\mathrm{\boldsymbol{\xi}}}}T\left(\widetilde{\rho}(\boldsymbol{\xi}),\rho(\boldsymbol{\xi;\theta})\right)\pi(\boldsymbol{\xi})d\boldsymbol{\xi}$ (13)
@@ 4 Learning Latent Dynamics of LLNL Testbed’s QPUs
We apply the learning strategy to identify the latent dynamics of two quantum processing units (QPUs), Dev1 and Dev2, at LLNL’s Quantum Design and Integration Testbed (QuDIT). Dev1 is a 2D single transmon made of Tantalum on a sapphire substrate [45]. Dev2 is a single 3D transmon mounted in a high-purity aluminum resonator. Both QPUs are mounted at 10 mK in a dilution fridge. The parameters of the base models (transition frequencies and coherence times) were estimated using standard characterization protocols, namely $\omega_{01}$, $T_{1}$ and $T_{2}$ were estimated using Rabi, Ramsey and energy decay protocols. The estimated parameters are given in Tab. 1 and show that Dev1 exhibits much longer coherence times than Dev2. We apply our data-driven technique on both QPUs to investigate its performance on devices that exhibit different noise levels.
Table 1: Parameters of two QPUs at LLNL’s Quantum Design and Integration Testbed. $\omega_{01}$ indicates the qubit transition frequency between $|0\rangle$ and $|1\rangle$. $T_{1}$ was measured via standard energy-decay experiment and $T_{2}$ via Ramsey experiments.
[…]
@@
[…]
$\vec{\rho}=\vec{M}^{-1}\vec{p}\quad\text{with}\quad\vec{p}=\begin{pmatrix}1\\ 2P(X)-1\\ 2P(Y)-1\\ 2P(Z)-1\end{pmatrix},\quad\vec{M}=\begin{pmatrix}1&0&0&1\\ 0&-1&-1&0\\ 0&i&-i&0\\ -1&0&0&1\end{pmatrix}$ (14)
[…]
Figure 2: Time evolution of the expected energy of a sample experiment (top) and expected trace distance (bottom) on Dev1 (left) and Dev2 (right) compared to the underlying base models.
@@ 4.1 Learning Dynamics of QPU - Dev1
We first investigate our UDE approach for learning dynamics of a more stable (but still noisy) Dev1 QPU. To investigate the effects of size of the training dataset, we train the models with data over two different durations, $T_{Tr.}=10~\mu s$ and $T_{Tr.}=20~\mu s$; the remaining data over $T_{Tr.}<t\leq 50~\mu s$ is used for validation. Here we learn a single source term $\mathcal{S}(\rho)$ using training data from a total of five experiments (i.e. $|\mathrm{\boldsymbol{\xi}}|=5$ in Eq. 12), in an attempt to learn an operator that is control-independent. We refer to such an operator as an Experiment-Generalized operator. The results obtained using the structure preserving ansatz will be denoted by $\mathcal{SP}$ and those obtained using the affine and nonlinear models will be denoted by $\mathcal{N}_{A}$ and $\mathcal{N}_{N}$, respectively.
Figure 4 shows the probability distribution of the trace distance between the density matrices obtained from experiments and those obtained using the trained ansatze for different base models (top vs bottom row) and for the two different training time domains (left vs right column). Investigating the effects of the size of the training dataset on accuracy, the higher, concentrated densities for low trace distances show that the accuracy of all models improves when training over larger intervals, both for the LvN and the Lindblad base model, and across all learning models. However, we observe that the accuracy improvement of the structure preserving model is marginal, suggesting that a smaller number of samples are needed to uniquely define such a UDE model. Comparing the accuracy of the different models over the training set, we observe from Figure 4 that all UDE models exhibit improved performance over the base models. Particularly, the structure-preserving model outperforms both the affine and nonlinear neural network models in accuracy over the training set, a trend observed across both training time intervals. The structure preserving and affine models outperform the nonlinear model in accuracy over the validation set; we observe this trend across both base models. The nonlinear neural network model, however, fails to enhance the accuracy of the already accurate Lindblad base model and only improves accuracy over the validation set when trained over the inaccurate LvN base model and on a more extended training interval.
@@
(a) Base model: Lindblad
Figure 4: Probability distribution of the trace-distance using UDE models trained with data from all experiments on Dev1 over $10\mu s$ (left) and $20\mu s$ (right). Dashed lines represent the training data set while solid lines denote the validation data set.
Table 2 presents the mean and standard deviation of the trace distance across various base models, UDE models, and training durations. In all scenarios, both the structure preserving and affine neural network models consistently deliver more accurate predictions, as indicated by their low expected values, and demonstrate greater stability, as evident by their reduced variance, when compared to either of the base models. Given the relatively noise-robust nature of this QPU, the superior performance of the structure preserving and affine neural network models suggests that they are more applicable when learning the underlying quantum dynamical processes than the nonlinear model. It’s worth mentioning that an alternative nonlinear model, considering different network architecture, activation functions, etc., may exhibit stronger performance for similar systems.
@@
where both operators show a detuning perturbation of approximately 11 kHz. The off-diagonal perturbations corresponding to $\sigma_{x}$ and $\sigma_{y}$ are on the order of 0.2 KHz and 2 KHz, respectively. Note that the difference between the two learned Hamiltonian perturbations are of the order of 0.6 kHz in the diagonal and 0.08 kHz in the off-diagonals, which can be attributed to the stochasticity introduced by mini-batching during model training. The estimated decoherence times $\mathcal{T}:=\left\{\gamma_{1}^{-1},\gamma_{2}^{-1},\gamma_{3}^{-1}\right\}$ when using the LvN and Lindblad base models are $\mathcal{T}^{LvN}=\left\{366,366,119\right\}~\mu s$ and $\mathcal{T}^{Lind}=\left\{1686,1686,688\right\}~\mu s$, respectively. Note that the magnitude of the dissipative perturbation are inversely proportional to the $\gamma_{j}^{-1}$. Hence, we see that the dissipative perturbation to the Lindblad base model is an order of magnitude smaller than it is to the LvN base model, as expected. Also, note that $\gamma_{1}=\gamma_{2}$ since $\widecheck{\Lambda}_{1}=i\widecheck{\Lambda}_{2}$ which corresponds to the collapse operator for energy decay. In all cases, we observed $\gamma_{1}=\gamma_{2}$, hence, here onwards, we report only on $\gamma_{1}$ and $\gamma_{3}$. The $T_{1}$ decay and $T_{2}$ dephasing times learned by the structure preserving ansatz for the LvN base model are $183~\mu s$ and $29.8~\mu s$, respectively. The effective $T_{1}$ decay and $T_{2}$ dephasing times (i.e. perturbed times) using the Lindblad base model are $171~\mu s$ and $27~\mu s$, respectively. The estimated $T_{2}$ dephasing times are comparable to those estimates using standard characterization protocols whereas the estimated $T_{1}$ decay times are significantly lower.
@@
As previously mentioned, the results depicted in Fig. 4 were obtained by learning a single operator that minimizes the expectation of the cost function over all experiments; we will refer to such operators as Experiment-Generalized (Exp-Gen) operators. However, in cases where the noise structure is itself experiment dependent, such an operator may not accurately capture the dynamics unique to a specific experiment. Hence, we attempt to learn operators tailored for individual experiments (i.e. $|\mathrm{\boldsymbol{\xi}}|=1$ in Eq. 12); we refer to such operators as Experiment-Specific (Exp-Spec) operators. Figure 5 presents the probability density of the trace distance over a single sample using different UDE models trained using Exp-Gen and Exp-Spec approaches. It is clear that all models outperform the two base models in accuracy. We observe that the Experiment-Specific operators yield more accurate predictions than the Experiment-General counterparts, with the nonlinear model demonstrating the largest improvement. The improvement in the Exp-Spec structure preserving model over the Exp-Gen model was largest when trained using the Lindblad base model but marginal when utilizing the LvN base model. The accuracy of the affine model using either base models was comparable.
@@
Figure 5: Probability distribution of the trace-distance using UDE models trained over a data from a single sample on Dev1 over a $20\mu s$ window. Dashed and solid lines represent Experiment-Generalized and Experiment-Specific operators, respectively.
@@
where we see that both base models converge to the same Hermitian perturbations. Furthermore, the perturbations are of a similar magnitude to those of the Exp-Gen. operators. The estimated perturbations to decoherence times when using the LvN and Lindblad base models are $\mathcal{T}^{LvN}=\left\{246,141\right\}~\mu s$ and $\mathcal{T}^{Lind}=\left\{580,1507\right\}~\mu s$, respectively. The effective $T_{1}$ decay times for both base models is $123~\mu s$. The $T_{2}$ dephasing times using the LvN and Lindblad base models are $35.2~\mu s$ and $29.4~\mu s$, respectively. Here, we see that despite using different base models, the structure preserving models converged to the same Hamiltonian and similar dissipative operators.

### v2

@@ 3.3 Training Procedure
The training procedure for learning the structure preserving model as well as the neural network-based source term amounts to solving an optimization that is constrained by the augmented dynamical equation (4). Denote by $\boldsymbol{\xi}$, a set of experiments that are parameterized by control functions which determine the time-evolution of the true/exact quantum state. For the examples below, this set consists of various microwave control pulses, $p(t),q(t)$ in (2) which will be described in Section 4. Each experiment defines thetrue/exact evolution of the density matrix, $\widetilde{\rho}(t_{j};\xi_{i})$, which is estimated using quantum state tomography [43] at various time steps $t_{j}\in(0,T]$. In this work, we apply linear inversion followed by spectral filtering and renormalization, Eq. 11. Let $\boldsymbol{\theta}$ denote the set of learnable parameters in the augmented differential equation, i.e. $\boldsymbol{\theta}=\left\{\alpha_{l}\bigcup\gamma_{l}:l=1,\ldots,N^{2}-1\right\}$ for the structure preserving ansatz in Section 3.1 and $\boldsymbol{\theta}=\left\{\mathbf{W}_{l}\bigcup\mathbf{b}_{l}:l=1,\ldots,L\right\}$ for the neural network model defined in Section 3.2. The goal of the training procedure is then to find optimal parameters $\boldsymbol{\theta}$ such that the resulting augmented differential equation matches the measured data $\widetilde{\rho}$ at time-steps $t_{j}$. To this end, the following constraint optimization problem is solved:
@@
In order to evaluate the objective function, the augmented dynamical system (12b) is evolved numerically for the given experimental setup in $\boldsymbol{\xi}$ and the current parameters $\mathrm{\boldsymbol{\theta}}$. Although any other numerical solver can be used, we here use an explicit 4th-order Runge-Kutta scheme for the numerical time-integration of the dynamical system. This scheme yields the predicted evolution $\left\{\rho(t_{j})\right\}_{j=1}^{N_{T}}$ which is compared to the true evolution within the loss function in terms of their Frobenius distance. We utilize the Julia SciML [44] framework to solve this optimization problem with the ADAM optimizer mini-batches as an initial optimizer, followed by a full-batch L-BFGS optimization. The gradient is calculated using automatic differentiation.
After training on a defined set of experiments and time-horizon $0<t_{j}<T$, we investigate the efficacy of the trained augmented model on a set of validation experiments. In particular, we show that the trained model is able to accurately predict the time-evolution on time domains much longer than the training domain, for experimental setups that have not been included in training. We refer to the accuracy over the training and validation sets as interpolation and extrapolation accuracy, respectively.
In the sections that follow, we will compare the solutions obtained using different formulations using the trace distance $T(\widetilde{\rho},\rho)=\frac{1}{2}\mathrm{Tr}\left[\sqrt{(\widetilde{\rho}-\rho)^{\dagger}(\widetilde{\rho}-\rho)}\right]$. In each case, we will apply the spectral filter in Eq. 11 to obtain a valid density matrix. To quantify the generalization (i.e. extrapolation) properties of the trained models, we employ the expected trace distance
[…]
@@ 4 Learning Latent Dynamics of LLNL Testbed’s QPUs
We apply the learning strategy to identify the latent dynamics of two quantum processing units (QPUs), Dev1 and Dev2, at LLNL’s Quantum Design and Integration Testbed (QuDIT). Dev1 is a 2D single transmon made of Tantalum on a sapphire substrate [45]. Dev2 is a single 3D transmon mounted in a high-purity aluminum resonator. Both QPUs are mounted at 10 mK in a dilution fridge. The parameters of the base models, such as transition frequencies ($\omega_{01}$), energy decay times ($T_{1}$), and dephasing times ($T_{2}$) were estimated using standard characterization protocols, using standard Rabi, Ramsey, and energy decay measurements. The measured parameters are given in Tab. 1. We apply our data-driven technique on both QPUs to investigate its performance on devices that exhibit different noise levels.
Table 1: Parameters of two QPUs at LLNL’s Quantum Design and Integration Testbed. $\omega_{01}$ indicates the qubit transition frequency between $|0\rangle$ and $|1\rangle$. $T_{1}$ was measured via standard energy-decay experiment and $T_{2}$ via Ramsey experiments. Dev1 exhibits much longer coherence times than Dev2.
QPU $\omega_{01}$ (GHz) $T_{1}$ ($\upmu$s) $T_{2}$ ($\upmu$s)
@@
To generate the training and validation data sets, we apply constant square pulses in the rotating frame at $\omega_{01}$. The set of experiments $\mathrm{\boldsymbol{\xi}}$ consists of various pulse amplitudes $\xi_{j}\sim\mathcal{U}(0,p_{max})$ with a maximum amplitude of $p_{max}=3.47$ MHz for Dev1 and $p_{max}=1.25$ MHz for Dev2. The pulses drive the qubits between the ground and first excited state (see Fig. 1). The pulses are applied for a total duration of $T=50~{}\upmu$s and $T=30~{}\upmu$s on Dev1 and Dev2, respectively, with measurements taken every 444~{}4ns. A total of 5000 shots are performed and the measured states are classified using a Gaussian Mixture Model (GMM). The density matrices at each timestep are then estimated by quantum state tomography using linear inversion estimate (LIE) [43]. The resulting Bloch vector is projected onto $x$, $y$, and $z$ basis, which yields the population vector $\mathrm{\boldsymbol{p}}$. The vectorized density matrix, $\mathrm{vec(\rho)}$, can then be obtained as
$\mathrm{vec(\rho)}=\mathrm{\boldsymbol{M}}^{-1}\mathrm{\boldsymbol{p}}\quad\text{with}\quad\mathrm{\boldsymbol{p}}=\begin{pmatrix}1\\ 2P(x)-1\\ 2P(y)-1\\ 2P(z)-1\end{pmatrix},\quad\mathrm{\boldsymbol{M}}=\begin{pmatrix}1&0&0&1\\ 0&-1&-1&0\\ 0&i&-i&0\\ -1&0&0&1\end{pmatrix}$ (22)
Figure 1 shows the time-evolution of the expected energy of the density matrices obtained through LIE for each device for a sample experiment, as well as the predicted evolution from the underlying base models in terms of Lindblad’s and Liouville-von-Neumann (LvN) equation. For Dev1, we see a good agreement between the experimental data and the Lindblad model; the LvN model shows comparable accuracy for short durations ($t\leq 6~{}\upmu$s) before the dissipative effects begin to dominate. We employ both the Lindblad and LvN equations as the base models when learning this QPU’s dynamics. In the case of Dev2, we see that the Lindblad model fails to accurately predict the population evolution over time. This suggests that the Lindblad model described by the Hamiltonian in Eq. 2, and $T_{1}$ and $T_{2}$ decoherence operators, does not capture all of the processes present on the testbed. Therefore, we only consider the Lindblad model as the baseline model when learning Dev2’s dynamics.
Figure 1: Time evolution of the expected energy of a sample experiment (top) and expected trace distance (bottom) on Dev1 (left) and Dev2 (right) compared to the underlying base models. Black line represents experimentally measured data.
@@ 4.1 Learning Dynamics of QPU - Dev1
We first investigate our UDE approach for learning dynamics of Dev1 QPU that has longer coherence times. To investigate the effects of size of the training dataset, we train the models with data over two different durations, $T_{Tr.}=10~{}\upmu$s and $T_{Tr.}=20~{}\upmu$s; the remaining data over $T_{Tr.}<t\leq 50~{}\upmu$s is used for validation. Here we learn a single source term $\mathcal{S}(\rho)$ using training data from a total of five experiments (i.e. $|\mathrm{\boldsymbol{\xi}}|=5$ in Eq. 3.3), to learn an operator that is control-independent. We refer to such an operator as an Experiment-Generalized operator. The results obtained using the structure preserving ansatz will be denoted by $\mathcal{SP}$ and those obtained using the affine and nonlinear models will be denoted by $\mathcal{N}_{A}$ and $\mathcal{N}_{N}$, respectively.
Figure 2 shows the probability distribution of the trace distance between the density matrices obtained from experiments and those obtained using the trained ansatze for different base models (top vs bottom row) and for the two different training time domains (left vs right column). Comparing the accuracy due to size of the training dataset, the higher and more concentrated densities for low trace distances show that the accuracy of all models improves when trained over larger intervals, both for the LvN and the Lindblad base model, and across all learning models. However, we observe that the improved accuracy of the structure preserving model is marginal, suggesting that a smaller number of samples are needed to uniquely define such a UDE model. Comparing the accuracy of the different models over the training set, we observe from Figure 2 that all UDE models exhibit improved performance over the base models. Particularly, the structure-preserving model outperforms both the affine and nonlinear neural network models in accuracy over the training set, while the structure preserving and affine models outperform the nonlinear model in accuracy over the validation set; we observe this trend across both base models. The nonlinear neural network model, however, fails to enhance the accuracy of the already accurate Lindblad base model and only improves accuracy over the validation set when trained over the inaccurate LvN base model and on a more extended training interval.
@@
(b) Base model: Lindblad
Figure 2: Probability distribution of the trace-distance using UDE models trained with data from all experiments on Dev1 over $10\upmu$s (left) and $20\upmu$s (right). Dashed lines represent the training data set while solid lines denote the validation data set.
Table 2 presents the mean and standard deviation of the trace distance across various base models, UDE models, and training durations. In all cases, both the structure preserving and affine neural network models consistently deliver more accurate predictions, as indicated by their low expected values, and demonstrate greater stability, as evident by their reduced variance, when compared to either of the base models. Given the relatively noise-robust nature of this QPU, the superior performance of the structure preserving and affine neural network models suggests that they are more applicable when learning the underlying quantum dynamical processes than the nonlinear model. It is worth mentioning that an alternative nonlinear model, considering different network architecture, activation functions, etc., may exhibit stronger performance for similar systems.
@@
where both operators show a detuning perturbation of approximately 11 kHz. The off-diagonal perturbations corresponding to $\sigma_{x}$ and $\sigma_{y}$ are on the order of 0.2 KHz and 2 KHz, respectively. Note that the difference between the two learned Hamiltonian perturbations are of the order of 0.6 kHz in the diagonal and 0.08 kHz in the off-diagonals, which can be attributed to the stochasticity introduced by mini-batching during model training. The estimated decoherence times $\mathcal{T}:=\left\{\gamma_{1}^{-1},\gamma_{2}^{-1},\gamma_{3}^{-1}\right\}$ when using the LvN and Lindblad base models are $\mathcal{T}^{LvN}=\left\{366,366,119\right\}~{}\upmu$s and $\mathcal{T}^{Lind}=\left\{1686,1686,688\right\}~{}\upmu$s, respectively. Note that the magnitude of the dissipative perturbation are inversely proportional to the $\gamma_{j}^{-1}$. Hence, we see that the dissipative perturbation to the Lindblad base model is an order of magnitude smaller than it is to the LvN base model, as expected. Also, note that $\gamma_{1}=\gamma_{2}$ since $\widecheck{\Lambda}_{1}=i\widecheck{\Lambda}_{2}$ which corresponds to the collapse operator for energy decay. In all cases, we observed $\gamma_{1}=\gamma_{2}$, hence, here onwards, we report only on $\gamma_{1}$ and $\gamma_{3}$. The $T_{1}$ decay and $T_{2}$ dephasing times learned by the structure preserving ansatz for the LvN base model are $183~{}\upmu$s and $29.8~{}\upmu$s, respectively. The effective $T_{1}$ decay and $T_{2}$ dephasing times (i.e. perturbed times) using the Lindblad base model are $171~{}\upmu$s and $27~{}\upmu$s, respectively. The estimated $T_{2}$ dephasing times are comparable to those estimates using standard characterization protocols whereas the estimated $T_{1}$ decay times are significantly lower.
@@
As previously mentioned, the results depicted in Fig. 2 were obtained by learning a single operator that minimizes the expectation of the cost function over all experiments; we will refer to such operators as Experiment-Generalized (Exp-Gen) operators. However, in cases where the noise structure is itself experiment dependent, such an operator may not accurately capture the dynamics unique to a specific experiment. Hence, we attempt to learn operators tailored for individual experiments (i.e. $|\mathrm{\boldsymbol{\xi}}|=1$ in Eq. 3.3); we refer to such operators as Experiment-Specific (Exp-Spec) operators. Figure 3 presents the probability density of the trace distance over a single sample using different UDE models trained using Exp-Gen and Exp-Spec approaches. It is clear that all models outperform the two base models in accuracy. We observe that the Exp-Spec operators yield more accurate predictions than the Exp-Gen counterparts, with the nonlinear model demonstrating the largest improvement. The improvement in the Exp-Spec structure preserving model over the Exp-Gen model was greatest when trained using the Lindblad base model but marginal when utilizing the LvN base model. The accuracy of the affine model using either base models was comparable.
@@
Figure 3: Probability distribution of the trace-distance using UDE models trained over a data from a single sample on Dev1 over a $20\upmu$s window. Dashed and solid lines represent Experiment-Generalized and Experiment-Specific operators, respectively.
@@
where both base models converge to the same Hermitian perturbations. Furthermore, the perturbations are of a similar magnitude to those of the Exp-Gen operators. The estimated perturbations to decoherence times when using the LvN and Lindblad base models are $\mathcal{T}^{LvN}=\left\{246,141\right\}~{}\upmu$s and $\mathcal{T}^{Lind}=\left\{580,1507\right\}~{}\upmu$s, respectively. The effective $T_{1}$ decay times for both base models is $123~{}\upmu$s. The $T_{2}$ dephasing times using the LvN and Lindblad base models are $35.2~{}\upmu$s and $29.4~{}\upmu$s, respectively. Here, despite using different base models, the structure preserving models converged to the same Hamiltonian and similar dissipative operators.